OpenStudy (anonymous):

Jeff bought 3 bags of dog food. He also bought a set of new dog bowls that cost eighteen dollars . He spent a total of forty-five dollars. Write both the verbal and mathematical equation to find out how much Jeff paid for each bag of dog food.

OpenStudy (anonymous):

Verbal model: Bags of dog food x cost of each bag cost of dog bowls = total cost. The equation is 3x 18 = 45. First, you subtract 18 from both sides. 45 - 18 = 27, and 18 - 18 is 0. Now the equation is 3x 0 = 27, so now we add 3x 0 = 3x. The equation is 3x = 27 now. Last, we divide both sides by 3. 27 / 3 = 9, and 3 / 3 = 1. So Jeff paid $9 for each bag of dog food.

OpenStudy (anonymous):

I did this before. I just copied my work and pasted here. :)

OpenStudy (anonymous):

WAIT COPY THIS ONE! Verbal model: Bags of dog food x cost of each bag cost of dog bowls = total cost. The equation is 3x + 18 = 45. First, you subtract 18 from both sides. 45 - 18 = 27, and 18 - 18 is 0. Now the equation is 3x 0 = 27, so now we add 3x 0 = 3x. The equation is 3x = 27 now. Last, we divide both sides by 3. 27 / 3 = 9, and 3 / 3 = 1. So Jeff paid $9 for each bag of dog food.

OpenStudy (anonymous):

what part